# Abelian congruences and similarity in varieties with a weak difference term

### v2

@@ Abstract.
This is the first of three papers motivated by the author’s desire to understand and explain “algebraically” one aspect of Dmitriy Zhuk’s proof of the CSP Dichotomy Theorem. In this paper we study abelian congruences in varieties having a weak difference term. Each class of the congruence supports an abelian group structure; if the congruence is minimal, each class supports the structure of a vector space over a division ring determined by the congruence. A construction due to J. Hagemann, C. Herrmann and R. Freese in the congruence modular setting extends to varieties with a weak difference term, and provides a “universal domain” for the abelian groups or vector spaces that arise from the classes of the congruence within a single class of the annihilator of the congruence. The construction also supports an extension of Freese’s similarity relation (between subdirectly irreducible algebras) from the congruence modular setting to varieties with a weak difference term.
@@ 2. Definitions
Let $\theta,\delta$ be congruences of an algebra $\mathbf{A}$}. The annihilator of $\theta$ modulo $\delta$, denoted $(\delta:\theta)$, is the unique largest congruence $\varphi$ for which $C(\varphi,\theta;\delta)$ holds. We say that $\theta$ is abelian if $C(\theta,\theta;0)$ holds; equivalently, if $\theta\leq(0:\theta)$. More generally, if $\theta,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\delta\leq\theta$, then we say that $\theta$ is abelian modulo $\delta$ if any of the following equivalent conditions hold: (i) $\theta/\delta$ is an abelian congruence of ${\mathbf{A}}/\delta$; (ii) $C(\theta,\theta;\delta)$ holds; (iii) $\theta\leq(\delta:\theta)$. The following lemma expresses that annihilators behave nicely under certain quotients and surjective homomorphisms.
@@ 3. Tools
The next lemma will be needed in Section 7. Its proof, in the congruence modular case, essentially can be extracted from the proof of [6, Theorem 10.11].
@@ 6.1. Easy case
(2) Let $\lambda,\mu\in F$. Observe that for each $\varphi$-class $C$, if $e$ is the unique element of $C\cap T$, then $\lambda{\upharpoonright}_{C}$ and $\mu{\upharpoonright}_{C}$ are endomorphisms of the abelian group $\operatorname{Grp}_{{\mathbf{D}}}(\varphi,e)$, $0{\upharpoonright}_{C}$ is the constant zero endomorphism of this group, and $(\lambda+\mu){\upharpoonright}_{C}$ is the pointwise sum of $\lambda{\upharpoonright}_{C}$ and $\mu{\upharpoonright}_{C}$ in this group. Thus the map $\Phi:F\rightarrow\prod_{e\in T}\operatorname{End}(\operatorname{Grp}_{{\mathbf{D}}}(\varphi,e))$ given by $\Phi(\lambda)_{e}=\lambda{\upharpoonright}_{C}$ where $C=e/\varphi$, is a unital embedding of $(F,+,-,\circ,0,1)$ into the direct product of the endomorphism rings of the groups $\operatorname{Grp}_{{\mathbf{D}}}(\varphi,e)$ for $e\in T$. Hence $(F,+,-,\circ,0,1)$ is a unital ring. It remains to show that each element of $F\setminus\{0\}$ is invertible.
Claim. $T$ is a maximal proper subuniverse of $\mathbf{D}$}.
Clearly $T\neq D$ since $\varphi\neq 0$ and $T$ is a transversal for $\varphi$. To prove the Claim, suppose $T<S\leq{\mathbf{D}}$ and $b\in D$; we will show $b\in S$. Fix $a\in S\setminus T$. Let $e=0(a)$ and $e^{\prime}=0(b)$, so $e,e^{\prime}\in T$ and $(a,e),(b,e^{\prime})\in\varphi$. Observe that $a\neq e$ (since $a\not\in T$), so $\operatorname{Cg}(a,e)=\varphi$, so $(b,e^{\prime})\in\operatorname{Cg}(a,e)$. By Corollary 3.2, there exists $f\in\operatorname{Pol}_{1}({\mathbf{D}})$ with $f(a)=b$ and $f(e)=e^{\prime}$. Pick a term $t(x,\mathbf{y})$ and parameters $c_{1},\ldots,c_{n}\in D$ with $f(x)=t(x,\mathbf{c})$. For each $i\in[n]$ let $e_{i}=0(c_{i})$, so $e_{i}\in T$ and $(c_{i},e_{i})\in\varphi$. Then
$t(e,e_{1},\ldots,e_{n})\stackrel{{\scriptstyle\varphi}}{{\equiv}}t(e,c_{1},\ldots,c_{n})=f(e)=e^{\prime}.$
Since $e^{\prime}$ and $t(e,e_{1},\ldots,e_{n})$ are in $T$ and $T$ is a transversal for $\varphi$, it follows that they are equal, so
(6.1) $t(\underline{e},e_{1},\ldots,e_{n})=t(\underline{e},c_{1},\ldots,c_{n}).$
As $\varphi$ is abelian, we can replace the underlined occurrences of $e$ in equation (6.1) with $a$ to get
(6.2) $t(a,e_{1},\ldots,e_{n})=t(a,c_{1},\ldots,c_{n}).$
Note that the left-hand side of equation 6.2 is in $S$, since $\{a,e_{1},\ldots,e_{n}\}\subseteq S$, while the right-hand side is $f(a)=b$, proving the Claim.
Now we finish the proof of (2). Let $\lambda\in F\setminus\{0\}$. One can easily check that $T\leq\operatorname{ran}(\lambda)\leq{\mathbf{D}}$ and $T\leq\lambda^{-1}(T)\leq{\mathbf{D}}$. Because $\lambda\neq 0$, we get $\operatorname{ran}(\lambda)\neq T$ and $\lambda^{-1}(T)\neq D$. Hence by the Claim, $\operatorname{ran}(\lambda)=D$ and $\lambda^{-1}(T)=T$. Thus $\lambda$ is surjective, and $\lambda{\upharpoonright}_{C}$ has trivial kernel for each $\varphi$-class $C$, so $\lambda$ is injective as well. Letting $\lambda^{-1}$ be the functional inverse of $\lambda$, we can easily check that $\lambda^{-1}\in F$. ∎
@@ 7. Similarity
In the special case where one of the two monoliths in the previous definition is nonabelian, the existence of a similarity bridge is possible only in the most trivial way.
@@
We end this section with two easy consequences of similarity in the abelian monolith case.
Proposition 7.9.
Suppose $\mathcal{V}$ is a variety with a weak difference term and ${\mathbf{A}},{\mathbf{B}}\in\mathcal{V}$ are subdirectly irreducible algebras with abelian monoliths $\mu,\kappa$ respectively. Define $\alpha=(0:\mu)$ and $\beta=(0:\kappa)$, and let $\mathbb{F}_{\mu}$ and $\mathbb{F}_{\kappa}$ be the division rings associated to $({\mathbf{A}},\mu)$ and $({\mathbf{B}},\kappa)$ respectively. If ${\mathbf{A}}\sim{\mathbf{B}}$, then:
  1. (1)
    ${\mathbf{A}}/\alpha\cong{\mathbf{B}}/\beta$.
  2. (2)
    $\mathbb{F}_{\mu}\cong\mathbb{F}_{\kappa}$.
Proof.
(1) Let $\lambda:D({\mathbf{A}})\cong D({\mathbf{B}})$. Let $\varphi_{{\mathbf{A}}},\varphi_{{\mathbf{B}}}$ be the monoliths of $D({\mathbf{A}}),D({\mathbf{B}})$ respectively. Then $\lambda$ induces an isomorphism $\lambda^{\ast}:D({\mathbf{A}})/\varphi_{{\mathbf{A}}}\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$. We have ${\mathbf{A}}/\alpha\cong D({\mathbf{A}})/\varphi_{{\mathbf{A}}}$ and ${\mathbf{B}}/\beta\cong D({\mathbf{B}})/\varphi_{{\mathbf{B}}}$ by Corollary 4.8, so ${\mathbf{A}}/\alpha\cong{\mathbf{B}}/\beta$.
(2) follows from Proposition 6.9. ∎

### v3

@@ Abstract.
This is the first of three papers motivated by the author’s desire to understand and explain “algebraically” one aspect of Dmitriy Zhuk’s proof of the CSP Dichotomy Theorem. In this paper we study abelian congruences in varieties having a weak difference term. Each class of the congruence supports an abelian group structure; if the congruence is minimal, each class supports the structure of a vector space over a division ring determined by the congruence. A construction due to J. Hagemann, C. Herrmann and R. Freese in the congruence modular setting extends to varieties with a weak difference term, and provides a “universal domain” for the abelian groups or vector spaces that arise from the classes of the congruence within a single class of the centralizer of the congruence. The construction also supports an extension of Freese’s similarity relation (between subdirectly irreducible algebras) from the congruence modular setting to varieties with a weak difference term.
@@ 2. Definitions
Let $\theta,\delta$ be congruences of an algebra $\mathbf{A}$}. The centralizer (or annihilator) of $\theta$ modulo $\delta$, denoted $(\delta:\theta)$, is the unique largest congruence $\varphi$ for which $C(\varphi,\theta;\delta)$ holds. We say that $\theta$ is abelian if $C(\theta,\theta;0)$ holds; equivalently, if $\theta\leq(0:\theta)$. More generally, if $\theta,\delta\in\operatorname{Con}({\mathbf{A}})$ with $\delta\leq\theta$, then we say that $\theta$ is abelian modulo $\delta$ if any of the following equivalent conditions hold: (i) $\theta/\delta$ is an abelian congruence of ${\mathbf{A}}/\delta$; (ii) $C(\theta,\theta;\delta)$ holds; (iii) $\theta\leq(\delta:\theta)$. The following lemma expresses that centralizers behave nicely under certain quotients and surjective homomorphisms.
@@ 3. Tools
The next two lemmas will be used in Sections 6. The proof of the second lemma, in the congruence modular case, can be extracted from the proof of [6, Theorem 10.11].
Lemma 3.10.
Suppose $\mathbf{A}$} is an algebra having a weak difference term, $\mu\in\operatorname{Con}({\mathbf{A}})$ is an abelian minimal congruence, and $S\leq{\mathbf{A}}$ is a subuniverse of $\mathbf{A}$} which is a transversal for $\mu$. Then $S$ is a maximal proper subuniverse of $\mathbf{A}$}.
Proof.
Clearly $S\neq A$ as $\mu\neq 0_{A}$. Let $\pi:{\mathbf{A}}\rightarrow{\mathbf{S}}$ be the homomorphism given by letting $\pi(x)$ be the unique element $y\in S$ satisfying $(x,y)\in\mu$. Fix $a\in A\setminus S$. Let $\rho=\operatorname{Sg}^{{\mathbf{A}}^{2}}(\{(a,\pi(a))\}\cup 0_{A})$. $\rho$ is a reflexive subuniverse of ${\mathbf{A}}^{2}$ satisfying $0_{A}\neq\rho\subseteq\mu$. Thus by Lemma 3.1 and minimality of $\mu$, we get $\rho=\mu$. Now let $b\in A$ be arbitrary. Then $(b,\pi(b))\in\mu=\rho$, so we can pick a term $t(x,\mathbf{y})$ and a tuple $\mathbf{c}=(c_{1},\ldots,c_{n})$ of elements of $A$ such that
$(b,\pi(b))=t^{{\mathbf{A}}^{2}}((a,\pi(a)),(c_{1},c_{1}),\ldots,(c_{n},c_{n}))=(t(a,\mathbf{c}),t(\pi(a),\mathbf{c})).$
Let $\mathbf{u}$ be the tuple in $S$ obtained by applying $\pi$ coordinatewise to $\mathbf{c}$. Observe that $\pi(b)=\pi(t(a,\mathbf{c}))=t(\pi(a),\mathbf{u})$, so
$t(\underline{\pi(a)},\mathbf{c})=t(\underline{\pi(a)},\mathbf{u}).$
As $\mathbf{c},\mathbf{u}$ are coordinatewise in $\mu$, $(a,\pi(a))\in\mu$, and $\mu$ is abelian, we can use Definition 2.2(3) and replace the underlined instances of $\pi(a)$ with $a$ to get $t(a,\mathbf{c})=t(a,\mathbf{u})$, proving $b=t(a,\mathbf{u})\in\operatorname{Sg}(S\cup\{a\})$. As $b$ was arbitrary, we get $\operatorname{Sg}(S\cup\{a\})=A$. ∎
@@ 6.1. Easy case
(2) Let $\lambda,\mu\in F$. Observe that for each $\varphi$-class $C$, if $e$ is the unique element of $C\cap T$, then $\lambda{\upharpoonright}_{C}$ and $\mu{\upharpoonright}_{C}$ are endomorphisms of the abelian group $\operatorname{Grp}_{{\mathbf{D}}}(\varphi,e)$, $0{\upharpoonright}_{C}$ is the constant zero endomorphism of this group, and $(\lambda+\mu){\upharpoonright}_{C}$ is the pointwise sum of $\lambda{\upharpoonright}_{C}$ and $\mu{\upharpoonright}_{C}$ in this group. Thus the map $\Phi:F\rightarrow\prod_{e\in T}\operatorname{End}(\operatorname{Grp}_{{\mathbf{D}}}(\varphi,e))$ given by $\Phi(\lambda)_{e}=\lambda{\upharpoonright}_{C}$ where $C=e/\varphi$, is a unital embedding of $(F,+,-,\circ,0,1)$ into the direct product of the endomorphism rings of the groups $\operatorname{Grp}_{{\mathbf{D}}}(\varphi,e)$ for $e\in T$. Hence $(F,+,-,\circ,0,1)$ is a unital ring.
It remains to show that each element of $F\setminus\{0\}$ is invertible. Let $\lambda\in F\setminus\{0\}$. One can easily check that $T\leq\operatorname{ran}(\lambda)\leq{\mathbf{D}}$ and $T\leq\lambda^{-1}(T)\leq{\mathbf{D}}$. Observe that $T$ is a maximal proper subuniverse of $\mathbf{D}$} by Lemma 3.10. Because $\lambda\neq 0$, we get $\operatorname{ran}(\lambda)\neq T$ and $\lambda^{-1}(T)\neq D$. Hence $\operatorname{ran}(\lambda)=D$ and $\lambda^{-1}(T)=T$. Thus $\lambda$ is surjective, and $\lambda{\upharpoonright}_{C}$ has trivial kernel for each $\varphi$-class $C$, so $\lambda$ is injective as well. Letting $\lambda^{-1}$ be the functional inverse of $\lambda$, we can easily check that $\lambda^{-1}\in F$. ∎
@@ 7. Similarity
Note also that, by Theorem 6.10, in case the monoliths $\mu$ and $\kappa$ are abelian, if ${\mathbf{A}}\sim{\mathbf{B}}$, then ${\mathbf{A}}/(0:\mu)\cong{\mathbf{B}}/(0:\kappa)$ and $\mathbb{F}_{\mu}\cong\mathbb{F}_{\kappa}$.
@@
In the special case where one of the two monoliths is nonabelian, the existence of a similarity bridge is possible only in the most trivial way.